OpenStudy (anonymous):

Fifty students in the third grade class listed their hair and eye colors in the table below:

OpenStudy (anonymous):

Brown hair Blonde hair Total Blue eyes 14 8 22 Brown eyes 16 12 28 39 20 50 Are the events "blonde hair" and "blue eyes" independent? Yes, P(blonde hair) • P(blue eyes) = P(blonde hair ∩ blue eyes) Yes, P(blonde hair) • P(blue eyes) ≠ P(blonde hair ∩ blue eyes) No, P(blonde hair) • P(blue eyes) = P(blonde hair ∩ blue eyes) No, P(blonde hair) • P(blue eyes) ≠ P(blonde hair ∩ blue eyes)

OpenStudy (anonymous):

@mathmale !

Miracrown (miracrown):

We need to check if that equation holds or not - so we need to find a few things like P(blonde hair) How many people are there in total?

OpenStudy (anonymous):

50?

Miracrown (miracrown):

Right! And how many of those have blonde hair?

OpenStudy (anonymous):

20

Miracrown (miracrown):

right! So... P(blonde hair) = 20/50 P(blue eyes) = 22/50 Agree? Now we are trying to find P(blonde hair AND blue eyes) right now that is where we are at.

Miracrown (miracrown):

p(Blonde hair and Blue eyes) = ?

OpenStudy (anonymous):

Would that be 8?

Miracrown (miracrown):

Yup, so it would be 8/50. :)

Miracrown (miracrown):

so now you have all these probabilities so now you can check if the equation holds

Miracrown (miracrown):

plug your numbers into that equation and see if that equation holds please

OpenStudy (anonymous):

OK, but how do I plug the numbers in? And to what equation?

Miracrown (miracrown):

p(blonde hair)*p(blue eyes) = p(blonde hair and blue eyes) that is the equation you need to determine if it is true you know each of those probabilities just plug the values into the right spots and see if the equation holds.

Miracrown (miracrown):

These are the probabilities: P(blonde hair) = 20/50 P(blue eyes) = 22/50 P(blonde hair and blue eyes= 8/50

OpenStudy (anonymous):

Oh, if its independent or not? So since 20 and 22 don't add up to 8, then it is independent?

Miracrown (miracrown):

the equation require you to multiply everything out 20/50 * 22/50 does that equal 8/50 ? that is what you need to determine right now

Miracrown (miracrown):

so what does 20/50 * 22/50 equal?

OpenStudy (anonymous):

No, it doesn't. 440/50

Miracrown (miracrown):

the bottom should be 50*50 right? or 2500

Miracrown (miracrown):

so does 440/2500 equal 8/50 ?

OpenStudy (anonymous):

right, sorry! No, so what does that mean?

Miracrown (miracrown):

so hopefully you stick with your answer from before and tell me that p(blonde hair)*p(blue eyes) is not the same thing as p(blonde hair and blue eyes) the fractions end up giving you something different on both sides so the equation didn't hold so looking at your answer choices, it is either B and D

Miracrown (miracrown):

so does this mean these events are independent ? when that equal fails to hold, does that mean they are independent?

Miracrown (miracrown):

so the answer is no here, they are not independent -- D they aren't independent because that equation isn't correct. :)

OpenStudy (anonymous):

I would say no.

Miracrown (miracrown):

great! :)

OpenStudy (anonymous):

Oh great, thanks for your help:)

Miracrown (miracrown):

its fine